OpenStudy (anonymous):

Solve the following pair of equations for x and y : 3x+2y=9xy ; 9x+4y=21xy ; x, y ≠ 0.

hartnn (hartnn):

do you know the method of elimination ?

hartnn (hartnn):

multiply 1st equation by 3 and subtract it from 2nd equation, what do u get ?

OpenStudy (anonymous):

yup elimination method...

OpenStudy (anonymous):

10y = 48 xy?

hartnn (hartnn):

did you add, instead of subtracting ? :P

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

cut the 9's

hartnn (hartnn):

9x+6y=27xy subtract from 9x+4y=21xy so, whats 6y-4y =... ?

OpenStudy (anonymous):

i thnk we shud multiply by -3 and add

OpenStudy (anonymous):

2y

hartnn (hartnn):

multiply by +3 and subtract is exactly equivalent to multiply by -3 and add do what you are comfortable with :)

hartnn (hartnn):

you should get 2y=6xy in the end, ask me if you don't get that

OpenStudy (anonymous):

yup...............finally gt it..................thnx man!!

hartnn (hartnn):

welcome ^_^